A lamp rated 120 watts and 240 volts what is its resistance
1
Expert's answer
2020-01-28T13:19:35-0500

We can find the resistance of the lamp from the definition of the electric power:


"P = IV,"

here, "P" is the power, "I" is the current flowing through the lamp and "V" is the voltage across the lamp.

We can find the current flowing through the lamp from the Ohm's law:


"I = \dfrac{V}{R},"

here, "R" is the resistance of the lamp.

Substituting "I" into the formula for the electric power, we get:


"P = IV = \dfrac{V^2}{R}."

From this formula we can find the resistance of the lamp:


"R = \dfrac{V^2}{P} = \dfrac{(240V)^2}{120W} = 480 \Omega."

Answer:

"R = 480 \Omega."